A certain perfect gas is compressed reversible from 1bar, 17°c to a pressure of 5bar in a perfect thermally insulated cylinder, the final temperature being 77°c
The work done on the gas during the compression is 45kj/kg
Calculate CV, R and the molecular weight of the gas
1
Expert's answer
2020-06-16T09:27:34-0400

Solution.

"p_1=1bar;"

"p_2=5bar;"

"T_1=290K;"

"T_2=350K;"

"\dfrac{W}{m}=45000J/kg;"

"W=mC_V(T_2-T_1)\implies C_V=\dfrac{W}{m(T_2-T_1)};" =\dfrac{W}{

"C_V=\dfrac{45000J/kg}{60K}=750J/(kg\sdot K);"

"\dfrac{T_2}{T_1}=(\dfrac{p_2}{P_1})^{\dfrac{\gamma-1}{\gamma}};"

"\dfrac{350}{290}=5^{\dfrac{\gamma-1}{\gamma}};" "1.2=5^{\dfrac{\gamma-1}{\gamma}};\dfrac{\gamma-1}{\gamma}=0.115; \gamma=1.3;"

"R=\dfrac{R_0}{M}\implies M=\dfrac{R_0}{R};"

"C_V=\dfrac{R}{\gamma-1}\implies R=C_V(\gamma-1);"

"R=750J/(kg\sdot K)(1.3-1)=225J/(kg\sdot K);"

"M=\dfrac{8.31}{225}=37\sdot10^{-3}kg/mol;"

Answer:"C_V=750J/(kg\sdot K);"

"R=225J/(kg\sdot K)" ;

"M=37\sdot10^{-3}kg/mol" .